# VideoCAD: A Large-Scale Video Dataset for Learning UI Interactions and 3D Reasoning from CAD Software

### v1

@@ 3.3 Benchmarking
Engineering UI is often more complex than traditional UI due to the necessity of many operations for precision control of the user. This can be seen in CAD software due to the complexity of the UI needed for precise spatial reasoning. To show this complexity, we benchmark VideoCAD extensively against other existing UI-agent datasets from literature in Table 1 (Appendix D).
@@
OSWorld 369 $15^{*}$
Mind2Web 2,350 7.3 1,135
WebArena 812
VisualWebArena 910 $35^{*}$
TheAgentCompany 175 40
WorkArena 33 15
WebShop 12,000 11.3 38
OmniAct 9,802
WebLinx 2,337 43 1,849
AITW 715,142 6.5
MMINA 1,050 12.9 601
MetaGUI 1,125 79
PixelHelp 187 4.2
AndroidWorld 116 18.1
AgentStudio 304 $30^{*}$
MoTIF 4,707 4.4 188
AndroidArena 116 11.4
WindowsAgentArena 154 8.1
MiniWoB++ 125 3.6 28
GUI-WORLD 12,379 10.97
@@
Table 1: Full Comparison of multi-environment benchmarks for GUI interaction.
The max is used instead of the average as the average is not reported
@@
As shown in Table 1, VideoCAD stands out across multiple dimensions. It features the longest time horizon—4$\times$ that of the next closest dataset (WebLinx)—and is the second-largest dataset after Android in the Wild, with over 50$\times$ more samples than the median (812). It is one of only two datasets requiring 3D spatial reasoning and pixel-level xy interactions, challenging agents to operate beyond text-based commands. For datasets with DOM access, VideoCAD also contains 6$\times$ more UI elements than the average web page in Mind2Web.
@@ 4 VideoCADFormer: An Autoregressive Transformer to Predict CAD Actions
We model CAD construction as a sequential decision process, where an agent observes UI frames and predicts low-level actions to recreate a target 3D model. Each trajectory in VideoCAD is a sequence $\tau=\{(\mathbf{I},\mathbf{o}_{t},\mathbf{a}_{t})\}_{t=0}^{T}$, where $\mathbf{I}$ is a fixed isometric view of the target shape, $\mathbf{o}_{t}$ is the UI frame at timestep $t$, and $\mathbf{a}_{t}$ is the expert action. The agent learns a policy $\pi_{\theta}(\mathbf{a}_{t}\mid\mathbf{o}_{t-k:t},\mathbf{I})$ via behavior cloning, trained to minimize a supervised loss over expert demonstrations $\mathcal{L}_{\text{BC}}=\mathbb{E}_{\tau\sim\mathcal{D}}\sum_{t}\ell\left(\pi_{\theta}(\mathbf{o}_{t-k:t},\mathbf{I}),\mathbf{a}_{t}\right),$ where $\ell$ is a structured prediction loss over commands and parameters. Our setup can be extended to reinforcement learning with rewards based on geometric similarity (see Appendix F for Chamfer Distance reward details).
@@
To interact with the CAD interface, the agent issues low-level UI commands. Each action $\mathbf{a}_{t}$ is a structured tuple: $\mathbf{a}_{t}=(c_{t},p_{t}^{1},\dots,p_{t}^{d_{t}})$, $c_{t}$ is the command type index, $d_{t}$ is the number of parameters for command $c_{t}$, $p_{t}^{i}$ is the $i$-th parameter for command $c_{t}$. Each action is a fixed-length vector: $\mathbf{a}_{t}=(c_{t},x_{t},y_{t},k_{t},n_{t},s_{t},v_{t})$, where each field corresponds to the parameter(s) used by one or more commands. Table 2 is the set of low-level UI commands and their associated parameters. Unused fields are padded with $-1$. All parameter values are discretized into 1000 classes, enabling the action prediction task to be cast as a multi-class classification problem. At the end of each CAD video, we set the part in isometric view; this action serves as our end-of-sequence command.
@@
Table 2: Structured action representation in VideoCAD. Each command type is mapped to a consistent 7D vector used for classification. Unused fields are set to $-1$.
@@ 4.1 Model Architecture
We propose VideoCADFormer, an autoregressive transformer tailored for CAD UI modeling (Figure 5). The model predicts low-level actions conditioned on a target CAD image, past UI-rendered frames, and prior actions. Visual inputs are processed by separate ViT encoders and fused via linear projection with timestep embeddings, capturing static goals and canvas evolution. Action tokens are embedded and passed to a causal transformer decoder with two key inductive biases: a causal mask to enforce autoregressive prediction, and a window mask to focus attention on recent context—reflecting the short-term dependency of UI commands. The decoder outputs command and parameter logits through two heads, with command-dependent masking to suppress irrelevant parameters. Full architectural details, training details and ablation studies are in Appendix E.
@@
We use two attention masks: a causal mask $\mathbf{M}^{\text{causal}}\in\mathbb{R}^{T\times T}$ is an upper-triangular matrix with $\infty$- ∞ in positions where future tokens should be masked and 00 elsewhere
@@
Command-dependent masks $M_{\hat{c}_{t}}\in\{0,1\}^{6}$ are applied to suppress unused parameter outputs. Invalid values are set to $-1$.
@@ 4.2 Evaluation Metrics
Geometric Fidelity. We evaluate geometric accuracy by executing model-predicted actions in Onshape and rendering the resulting CAD models. On 100 test sequences, we compute mean bidirectional Chamfer Distance (CD) after PCA alignment (Appendix F). A sequence is considered successful if CD $<$ 0.02, a threshold chosen empirically based on human evaluation, where shapes below this value are visually indistinguishable (Appendix F). Invalid sequences (i.e., those failing to mesh) are also reported.
@@ 5 Results
Accuracy and Perfect Sequences. We use Video Pre-training (VPT) baker2022videopretrainingvptlearning (a leading method in offline behavior cloning for Minecraft), Pix2Act shaw2023pixelsuiactionslearning and Pearce et. al. pearce2021counterstrikedeathmatchlargescalebehavioural as baselines. As shown in Table E, VideoCADFormer achieves the highest command and parameter accuracy across all evaluated models, outperforming VPT, Pix2Act, and Pearce et al. It also leads across all metrics of perfect action sequence prediction. The model shows consistent improvements over VPT across all levels of task difficulty. These results highlight VideoCADFormer’s robustness in executing complete CAD sequences without error, particularly in long-horizon settings where small mistakes compound. The increasing performance gap with task difficulty indicates better generalization under growing complexity.
Geometric Fidelity. Table 4 evaluates the final CAD quality by executing the predicted sequences in Onshape. VideoCADFormer outperforms VPT in both short and long sequences in terms of Chamfer-based success rate, while VPT performs slightly better on medium sequences. Overall, VideoCADFormer achieves lower Chamfer distance and a lower proportion of invalid CAD models. These results suggest that stronger sequence-level prediction directly translates to more accurate 3D model reconstruction.
@@
Table 4: Performance by sequence length on Chamfer success rate ($<$0.02 $\uparrow$), mean Chamfer distance ($\downarrow$), and invalid sample rate ($\downarrow$). Overall metrics are averaged across categories.
@@ 5.1 Goal-Driven CAD Generation and Autocompletion
Shape Generation from Scratch. Conditioned only on a target CAD image, our model can generate full action sequences that construct the corresponding geometry from an empty canvas. As shown in Figure 6, the predicted low-level UI actions reliably recreate complex multi-step CAD models. This demonstrates the model’s capacity to plan and execute long-horizon sequences that align with spatial goals purely from visual context.
@@
Failure Cases and Limitations. Model failures are primarily due to inaccurate $(x,y)$ predictions that cause open or slightly distorted sketch loops, preventing valid extrusions. Misclassification between lines and arcs also occurs, especially when curvature is ambiguous. These errors (Figure 11) are often minor and correctable, but they highlight the limitations of image-only supervision and suggest the need for topological constraints or interactive fine-tuning. More details in Appendix H.
@@ 5.2 Case Study: Evaluating LLMs on 3D Reasoning and CAD Understanding
LLMs Struggle with 3D Reasoning and CAD Video Understanding.
To evaluate the limits of modern multimodal LLMs in spatially grounded engineering domains, we construct a VQA benchmark, VideoCAD VQA, which is derived from VideoCAD. This benchmark contains multiple-choice multimodal questions that probe fine-grained understanding of both temporal video sequences and 3D geometric properties—tasks essential for CAD modeling. It includes four task types: action recognition, frame ordering, extrusion counting, and depth comparison. Questions are generated automatically from ground-truth UI logs and CAD geometry (Appendix I).
@@
[…]
google2025gemini25 random
Extrusion Shape Prediction 27.0 22.5 19.5 22.5 25.5 29.0
Number of Extrusion Estimation 47.0 37.5 47.0 45.0 38.0 21.4
Extrusion Difference Prediction 73.5 80.0 62.0 71.5 71.0 49.1
Sketch Ordering 37.5 29.5 41.0 37.0 61.0 35.0
Sketch Identification 62.0 48.5 43.5 48.5 59.0 21.5
Plane Identification 87.0 86.5 86.0 91.5 89.5 34.0
CAD Primitive Identification 84.1 80.1 78.6 75.1 82.6 40.3
Sequence Prediction 81.0 68.0 70.5 77.5 79.2 36.5
Video Frame Sequencing 36.0 23.0 32.5 80.0 73.2 17.4
Hole Detection 92.0 53.5 79.5 88.4 81.0 50.0
Symmetry Detection 18.5 19.0 12.0 27.9 27.0 12.5
Table 5: Performance of various vision-language models on the CAD-VQA benchmark across 11 evaluation tasks. Metrics are accuracy (%).
We perform 0-shot evaluation on VideoCAD VQA by feeding a template followed by the prompt to each LLM. If there is no valid answer in the model’s response, we perform random selection as a remedy. We evaluate the model 3 times for each question. Table LABEL:tab:cad_vqa shows results across multimodal LLMs. Despite their strong performance on general VQA benchmarks, current models underperform across most categories. For instance, GPT-4.1 achieves only 47% accuracy on extrusion number estimation and 36% on frame ordering, highlighting persistent challenges in grounded visual reasoning and temporal understanding in CAD workflows. These results underscore a critical gap in LLMs’ ability to interpret complex video-based modeling workflows. Importantly, this VQA benchmark serves as a case study showcasing how VideoCAD enables rigorous evaluation of spatial and procedural understanding in LLMs. Many additional multimodal tasks—such as symmetry detection, subpart recognition, or parameter estimation—can be derived from the same dataset and are discussed in Appendix I.
@@
We also benchmark LLMs as UI agents in complex engineering software. Using BrowserGym dechezelles2025browsergymecosystemwebagent , we prompt models to perform CAD construction tasks within Onshape based only on a target screenshot. Each agent is given 200 steps and must produce xy coordinates corresponding to UI actions (e.g., drawing, extrusion). Tasks are drawn from the 10 shortest sequences in VideoCAD, with early termination if no meaningful canvas modification occurs. No LLM—including GPT-4.1, Claude-3.7, and Gemini-2.5—is able to successfully complete a full CAD construction. We anticipate that this happens due to the need for pixel-level precision, long-horizon spatial planning, and consistent 3D reasoning capabilities for CAD modeling unlike web tasks (e.g., spreadsheets or browser automation  liang2024taskmatrix ).
@@ 6 Limitations and Future Work
While VideoCAD offers a high-fidelity benchmark for CAD UI modeling, it has several limitations. All trajectories are synthetically generated by a rule-based bot, and despite human-inspired heuristics, they lack natural variability in timing, errors, and strategy. The dataset focuses solely on sketch-extrude workflows, omitting advanced operations like fillets, sweeps, and lofts. Interactions are limited to a single platform—OnShape—raising concerns about generalization. Finally, our current end-to-end benchmarks are only validated on a small subset due to the high cost of CAD rendering and geometric comparison. To address these gaps, future work will: (1) incorporate human demonstration data (e.g., from YouTube CAD tutorials), (2) extend coverage to assemblies and more advanced CAD features, (3) support additional CAD platforms (e.g., Fusion 360, FreeCAD), (4) collect multiple trajectories per CAD target to capture variation across users, (5) introduce extrusion-by-extrusion text prompts to enable natural interaction between users and AI agents in CAD softwares.
@@ 7 Conclusion
Appendix A Onshape

### v2

@@ 3.3 Benchmarking
Engineering UI is often more complex than traditional UI due to the necessity of many operations for precision control of the user. This can be seen in CAD software due to the complexity of the UI needed for precise spatial reasoning. To show this complexity, we benchmark VideoCAD extensively against other existing UI-agent datasets from the literature in Table LABEL:tab:environment_comparison (Appendix LABEL:app:dataset-bench).
@@
OSWorld xie2024osworldbenchmarkingmultimodalagents 369 $15^{*}$
Mind2Web deng2023mind2webgeneralistagentweb 2,350 7.3 1,135
WebShop yao2023webshopscalablerealworldweb 12,000 11.3 38
WebLinx lù2024weblinx 2,337 43 1,849
AITW rawles2023androidwildlargescaledataset 715,142 6.5
MMINA zhang2024mminabenchmarkingmultihopmultimodal 1,050 12.9 601
MetaGUI sun2022metaguimultimodalconversationalagents 1,125 79
MoTIF burns2022motifvln 4,707 4.4 188
GUI-WORLD chen2025guiworldvideobenchmarkdataset 12,379 10.97
@@
Table 1: Comparison of multi-environment benchmarks for GUI interaction. The max is used instead of the average as the average is not reported.
@@
[…]
@@ 4 VideoCADFormer: An Autoregressive Transformer to Predict CAD Actions
We model CAD construction as a sequential decision process, where an agent observes UI frames and predicts low-level actions to recreate a target 3D model. Each trajectory in VideoCAD is a sequence $\tau=\{(\mathbf{I},\mathbf{o}_{t},\mathbf{a}_{t})\}_{t=0}^{T}$, where $\mathbf{I}$ is a fixed image of the target shape, $\mathbf{o}_{t}$ is the UI frame at timestep $t$, and $\mathbf{a}_{t}$ is the expert action. The agent learns a policy $\pi_{\theta}(\mathbf{a}_{t}\mid\mathbf{o}_{t-k:t},\mathbf{I})$ via behavior cloning, trained to minimize a supervised loss over expert demonstrations $\mathcal{L}_{\text{BC}}=\mathbb{E}_{\tau\sim\mathcal{D}}\sum_{t}\ell\left(\pi_{\theta}(\mathbf{o}_{t-k:t},\mathbf{I}),\mathbf{a}_{t}\right),$ where $\ell$ is a structured prediction loss over commands and parameters. Our setup can be extended to reinforcement learning with rewards based on geometric similarity (see Appendix C for Chamfer Distance reward details).
@@
To interact with the CAD interface, the agent issues low-level UI commands. Each action $\mathbf{a}_{t}$ is a structured tuple: $\mathbf{a}_{t}=(c_{t},p_{t}^{1},\dots,p_{t}^{d_{t}})$, $c_{t}$ is the command type index, $d_{t}$ is the number of parameters for command $c_{t}$, $p_{t}^{i}$ is the $i$-th parameter for command $c_{t}$. Each action is a fixed-length vector: $\mathbf{a}_{t}=(c_{t},x_{t},y_{t},k_{t},n_{t},s_{t},v_{t})$, where each field corresponds to the parameter(s) used by one or more commands. Table 2 is the set of low-level UI commands and their associated parameters. Unused fields are padded with 1-1. All parameter values are discretized into 1000 classes, enabling the action prediction task to be cast as a multi-class classification problem. At the end of each CAD video, we set the part in isometric view; this action serves as our end-of-sequence command.
@@
Table 2: Structured action representation in VideoCAD. Each command type is mapped to a consistent 7D vector used for classification. Unused fields are set to 1-1.
@@ 4.1 Model Architecture
We propose VideoCADFormer, an autoregressive transformer tailored for CAD UI modeling (Figure 4). The model predicts low-level actions conditioned on a target CAD image (sketch, kernel-rendered, or realistic), past UI-rendered frames, and prior actions. Visual inputs are processed by separate ViT encoders and fused via linear projection with timestep embeddings, capturing static goals and canvas evolution. Action tokens are embedded and passed to a causal transformer decoder with two key inductive biases: a causal mask to enforce autoregressive prediction, and a window mask to focus attention on recent context—reflecting the short-term dependency of UI commands. The decoder outputs command and parameter logits through two heads, with command-dependent masking to suppress irrelevant parameters. Full training details and ablation studies are in Appendix B.
@@
We use two attention masks: a causal mask $\mathbf{M}^{\text{causal}}\in\mathbb{R}^{T\times T}$ is an upper-triangular matrix with -$\infty$ in positions where future tokens should be masked and 0 elsewhere
@@
Command-dependent masks $M_{\hat{c}_{t}}\in\{0,1\}^{6}$ are applied to suppress unused parameter outputs. Invalid values are set to 1-1.
@@ 4.2 Evaluation Metrics
Geometric Fidelity. We evaluate geometric accuracy by executing model-predicted actions in Onshape and rendering the resulting CAD models. On 200 test sequences, we compute mean bidirectional Chamfer Distance (CD) after PCA alignment (Appendix C). A sequence is considered successful if CD << 0.02, a threshold chosen empirically based on human evaluation, where shapes below this value are visually indistinguishable (Appendix C). Invalid sequences (i.e., those failing to mesh) are also reported.
@@ 5 Results
Accuracy and Perfect Sequences. We use Video Pre-training (VPT) baker2022videopretrainingvptlearning (a leading method in offline behavior cloning for Minecraft), Pix2Act shaw2023pixelsuiactionslearning and Pearce et. al. pearce2021counterstrikedeathmatchlargescalebehavioural as baselines. As shown in Table B, VideoCADFormer achieves the highest command and parameter accuracy across all evaluated models, outperforming VPT, Pix2Act, and Pearce et al. It also leads across all metrics of perfect action sequence prediction. The model shows consistent improvements over VPT across all levels of task difficulty. These results highlight VideoCADFormer’s robustness in executing complete CAD sequences without error, particularly in long-horizon settings where small mistakes compound. The increasing performance gap with task difficulty indicates better generalization under growing complexity.
Geometric Fidelity. Table 4 evaluates the final CAD quality by executing the predicted sequences in Onshape. VideoCADFormer outperforms VPT in both short and long sequences in terms of Chamfer-based success rate, while VPT performs slightly better on medium sequences. Overall, VideoCADFormer achieves a lower Chamfer distance and a lower proportion of invalid CAD models. These results suggest that stronger sequence-level prediction directly translates to more accurate 3D model reconstruction.
@@
Table 4: Performance by sequence length on Chamfer success rate (<<0.02 $\uparrow$), mean Chamfer distance ($\downarrow$), and invalid sample rate ($\downarrow$). Overall metrics are averaged across categories.
@@ 5.1 Goal-Driven CAD Generation and Autocompletion
Shape Generation from Scratch. Conditioned only on a target CAD image, our model can generate full action sequences that construct the corresponding geometry from an empty canvas. As shown in Figure 5, the predicted low-level UI actions reliably recreate complex multi-step CAD models. This demonstrates the model’s capacity to plan and execute long-horizon sequences that align with spatial goals purely from visual context. Furthermore, we train models to take any of the three image types—sketch, kernel-rendered, or realistic—as input representations of the target CAD geometry (see Appendix B for details).
@@
Failure Cases and Limitations. Model failures are primarily due to inaccurate $(x,y)$ predictions that cause open or slightly distorted sketch loops, preventing valid extrusions. Misclassification between lines and arcs also occurs, especially when curvature is ambiguous. These errors (Figure 10) are often minor and correctable, but they highlight the limitations of image-only supervision and suggest the need for topological constraints or interactive fine-tuning. More details in Appendix F.
@@ 5.2 Case Study: Evaluating LLMs on 3D Reasoning and CAD Understanding
Comparison
gpt-4.1-2025-04-14 openai2025gpt41 84.1 36.0 47.0 73.5
claude-3-7-sonnet-20250219anthropic2025claude37 80.1 23.0 37.5 80.0
qwen2.5-vl-72b-instructqwen2025vl72b 78.6 32.5 47.0 62.0
o3-2025-04-16 openai2025o3 75.1 80.0 45.0 71.5
gemini-2.5-pro-preview-05-06 google2025gemini25 82.6 73.2 38.0 71.0
random 40.3 17.4 21.4 49.1
Table 5: Comparison of models on CAD reasoning tasks. Metrics are in %.
LLMs Struggle with 3D Reasoning and CAD Video Understanding.
A key capability for AI models to create 3D CAD designs is to excel in 3D reasoning. To evaluate the limits of modern multimodal LLMs in spatially grounded engineering domains, we construct a VQA benchmark, VideoCAD VQA, which is derived from VideoCAD. This benchmark contains multiple-choice multimodal questions that probe fine-grained understanding of both temporal video sequences and 3D geometric properties—tasks essential for CAD modeling. It includes four task types: action recognition, frame ordering, extrusion counting, and depth comparison. Questions are generated automatically from ground-truth UI logs and CAD geometry (Appendix H).
We perform zero-shot evaluation on VideoCAD VQA by feeding a template followed by the prompt to each LLM. If there is no valid answer in the model’s response, we perform random selection as a remedy. We evaluate the model three times for each question. Table 5 shows results across multimodal LLMs. Despite their strong performance on general VQA benchmarks, current models underperform across most categories. For instance, GPT-4.1 achieves only $47\%$ accuracy on extrusion counting and $18\%$ on depth estimation, which highlights persistent challenges in grounded visual reasoning and geometry-aware token alignment. These results underscore a critical gap in LLMs’ ability to interpret complex video-based modeling workflows. Importantly, this VQA benchmark serves as a case study showcasing how VideoCAD enables rigorous evaluation of spatial and procedural understanding in LLMs. Many additional multimodal tasks—such as symmetry detection, subpart recognition, or parameter estimation—can be derived from the same dataset and are discussed in Appendix H.
@@
We also benchmark LLMs as UI agents in complex engineering software. Using BrowserGym dechezelles2025browsergymecosystemwebagent, we prompt models to perform CAD construction tasks within Onshape based only on a target screenshot. Each agent is given 200 steps and must produce xy coordinates corresponding to UI actions (e.g., drawing, extrusion). Tasks are drawn from the 10 shortest sequences in VideoCAD, with early termination if no meaningful canvas modification occurs. All LLMs—including GPT-4.1, Claude-3.7, and Gemini-2.5—fail to complete a full CAD construction (Appendix ). We anticipate that this happens due to the need for pixel-level precision, long-horizon spatial planning, and consistent 3D reasoning capabilities for CAD modeling unlike web tasks (e.g., spreadsheets or browser automation  liang2024taskmatrix).
@@ 6 Limitations and Future Work
While VideoCAD offers a high-fidelity benchmark for CAD UI modeling, it has several limitations. All trajectories are synthetically generated by a rule-based bot, and despite human-inspired heuristics, they lack natural variability in timing, errors, and strategy. The dataset focuses solely on sketch-extrude operations, omitting operations like fillets, sweeps, and lofts. Interactions are limited to a single platform—Onshape—raising concerns about generalization to other CAD software. Finally, our current end-to-end benchmarks are only validated on a small subset due to the high cost of CAD rendering and geometric comparison. To address these gaps, future work will: (1) incorporate human demonstration data (e.g., from YouTube CAD tutorials), (2) extend coverage to assemblies and more advanced CAD features, (3) support additional CAD platforms (e.g., Fusion 360, FreeCAD), (4) collect multiple trajectories per CAD target to capture variation across users, (5) introduce extrusion-by-extrusion text prompts to enable natural interaction between users and AI agents in CAD softwares.
@@ 7 Conclusion
Table of Contents for Appendices
Appendix A CAD Software